# Constructing embedded surfaces for cellular embeddings of leveled spatial graphs

### v1

@@ Abstract
Finding a closed orientable surface $\mathcal{S}$ embedded in $\mathbb{R}^{3}$ where a given spatial graph $\mathcal{G}\subset\mathbb{R}^{3}$ cellular embeds is in general not possible. We therefore restrict our interest to the special class of spatial graphs that are leveled. We show that for leveled spatial graphs with a small number of levels, a surface $\mathcal{S}$ can always be found. The argument is based on the idea of decomposing $\mathcal{G}$ into subgraphs that can be placed on a sphere and on handles that are attached to the sphere, together forming an embedding of $\mathcal{G}$ in $\mathcal{S}$. We generalize the procedure to an algorithm that, if successful, constructs $\mathcal{S}$ for leveled spatial graphs with any number of levels. We conjecture that all connected leveled embeddings can be cellular embedded with the presented algorithm.
@@ 1 Introduction
Cellular embeddings, as a natural generalization of planar graphs, have in particular been studied in topological graph theory and combinatorics. They provide a combinatorial description of graph embeddings in surfaces [9], and they are at the core of long-standing open conjectures such as the circular (or strong) embedding conjecture [6].
@@
Given that graphs on surfaces have mainly be studies in topological graph theory that, in contrast to knot theory, traditionally considers all possible embeddings of an abstract graph, the scenario involving specific realizations of graphs within three-dimensional space, where surfaces must not intersect, remains underexplored. This is despite the setting being most relevant for applications taking place in the 3-dimensional Euclidean world. For example, building a graph-like structure using self-assembly of DNA molecules by origami folding requires first finding an unknotted A-trail through the desired spatial graph [4], [7]. When the graph is eulerian and cellular embedded in a surface, an unknotted A-trail can be found by following the boundary of the disks [4].
@@
Since we do not require the surface to be standardly embedded (e.g. a torus is allowed to be knotted), it is in principle possible to obtain all cellular embeddings of spatial graphs by starting from cellular embeddings of free bouquets of circles, on which subdivisions of edges and splittings of vertices are performed on the surface. However, our goal is to construct a surface for a given spatial graph. Note that rotation systems, although equivalent to cellular embeddings for abstract graphs [9], do not ensure the described surface to be free of self-intersections.
@@
A natural approach to construct a surface for a given spatial graph is to glue together non-intersecting properly embedded disks that are bounded by some cycles of the graph. A sufficient albeit unnecessarily strong condition for the existence of such disks is $k$-flatness of the spatial graph [8]. Indeed, it suffices to find a collection of cycles that cover all edges of the graph such that each edge is covered at most twice, and that bound disks satisfying the property above. The union of the disks and the spatial graph produces a cellular embedding of the spatial graph in the resulting surface. In case that every edge belongs to exactly two different cycles, the embedding is a circular (or strong) embedding. However, finding a collection of such disks is in general an intractable problem, as it includes detecting the unknot, non-trivial links with any number of components, Borromean links, and more complicated entanglements such as nonfree knot- and linkless spatial graphs such as the Kinoshita theta-curve.
@@
We therefore follow another approach to construct the surface. Without loss of generality, we can restrict to connected spatial graphs. Indeed, if distinct connected components are nontrivially linked, the spatial graph cannot be cellular embedded, and unlinked components can be considered independently. The idea is to start with a subgraph of the spatial graph that cellular embeds in the sphere, and then iteratively attach handles on which the remaining parts of the graph sit on. If the graph follows a longitude and a meridian of every handle, the resulting embedding is cellular. Since in our construction a longitude of the cylinder that generates a new handle is always followed by some subgraph, cellularity is implied if the two boundary components of the cylinder are attached to different disks of the already generated surface. If the two boundary components of the cylinder are attached to the same disk, it must be ensured that the graph also runs through the meridian of the resulting handle. Furthermore, conditions need to be found that guarantee that the construction results in an embedded surface, i.e., handles should be attached in such a way that they do not intersect the already constructed surfasce parts. For such a procedure to work, we require the spatial graph to have some properties and we call such spatial graphs leveled. They can be viewed as a generalization of planar graphs and are distinct from but related to open book embeddings [2].
@@ 2 Preliminaries
All graphs $G$ are undirected finite graphs. Graphs are allowed to have loops and parallel edges. The fragments (or bridges) of a graph $G$ with respect to a cycle $C$ of $G$ are the closures of the connected components of $G-C$ [10]. Two fragments are conflicting if they either have pairs of endpoints on $C$ that alternate, or have at least three endpoints in common [11]. The conflict graph of a graph $G$ with respect to a cycle $C$ is the graph that has a vertex for each fragment of $G$ with respect to $C$, and two vertices are adjacent if and only if the corresponding fragments conflict. A connected graph is hamiltonian if it contains a cycle that visits every vertex exactly once, i.e., all its fragments with respect to the cycle are the closures of edges or loops. A graph embedding is an embedding $f:G\rightarrow S^{3}$ of a graph $G$ in $S^{3}$ up to ambient isotopy. The corresponding spatial graph $\mathcal{G}$ is the image of this embedding. A spatial graph is trivial if it embeds in $S^{2}$. An abstract graph is planar if it admits a trivial embedding. A diagram of a spatial graph $\mathcal{G}$ is the image of a projection of $\mathcal{G}$ onto $\mathbb{R}^{2}$ such that all intersections in the image are transversal, no vertices and at most two points are mapped to a crossing, and each double point is assigned as an over- or under-crossing. We focus on a special class of spatial graphs, the leveled graphs:
@@
We define an ordering of the fragments of $\mathcal{G}$ as follows: Fix a diagram $D$ of a leveled spatial graph $\mathcal{G}$ such that the spine $C$ of $\mathcal{G}$ does not have crossings on $D$. If the fragment $f$ crosses over fragment $g$, then $f>g$. Fragments that are not comparable by the partial order are said to be on the same level. Note that conflicting fragments must be on different levels.
This definition allows to assign the fragments of a leveled spatial graph to their level. The fragments that do not cross over any other fragments are at level $1$, denoted $L_{1}$.
@@
If the leveled spatial graph $\mathcal{G}$ has $n$ levels $L_{i},1\leq i\leq n$, there is an action of the cyclic group $C_{n}=\langle x\mid x^{n}=1\rangle$ on the partition of levels =$\mathcal{L}=\{L_{1},\dots,L_{n}\}$. The action is given by sending $x\cdot L_{i}$ to $L_{i+1(\mathrm{mod}\ n)}.$
We say that a leveled spatial graph $\mathcal{G}$ is in representative leveling if for each fragment $f$ at level $i>1$, there exists a fragment $g$ at level $i-1$ such that $f$ and $g$ conflict. In other words, $\mathcal{G}$ is in representative leveling if every fragment belongs to the lowest possible level. Every leveled spatial graph can be put into representative leveling by assigning fragments to the lowest possible level. Note that the cyclic reordering of levels does not maintain representative leveling.
Every connected trivial spatial graph is a leveled graph with exactly one level. Consequently, all planar graphs admit a leveled embedding, but this is not true for all graphs. For example, the graph $K_{5}\_K_{5}$ that is obtained by joining two copies of $K_{5}$ by an edge (fig. 11) cannot be embedded as a leveled graph, since every choice of a cycle implies a nonplanar fragment. Therefore, no spine can be found for $K_{5}\_K_{5}$.
@@
If the two fragments that have the vertex $v$ in common are non-conflicting, following $C$ starting from $v$ runs through all endpoints of one fragment before running through all endpoints of the other. Splitting $v$ as described at above does not introduce an alternation of the endpoints, and therefore keeps the fragments non-conflicting.
By repeatedly splitting vertices on $C$, their degree can be reduced to three. ∎
A hamiltonian leveled spatial graph $\mathcal{H}$ is a leveled spatial graph whose spine is a hamiltonian cycle of the underlying abstract graph. For constructing a cellular embedding for a hamiltonian leveled spatial graph, due to lemma 1, it is possible to restrict without loss of generality to simple graphs, i.e. graphs without loops and parallel edges. In this case, all fragments are edges with two distinct endpoints. Hamiltonian leveled spatial graphs can be completely described combinatorially by their spine list that encodes the order in which fragments are attached to the spine together with their level. If several fragments share an endpoint, the order in the spine list is chosen such that the alternation of fragment labels is avoided. The spine list is unique up to permutation of the labels of parallel edges and cyclic permutations of its symbols, which corresponds to different choices of start points on the spine, see fig. 1.
@@ 3 Cellular embeddings of leveled spatial graphs with few levels
Assume $\mathcal{G}$ is in representative leveling. Let $L_{i}$ be the union of all fragments of level $i$, and let $L_{1,2}:=C\cup L_{1}\cup L_{2}$ be the subgraph of $\mathcal{G}$ that consists of the spine $C$ of $\mathcal{G}$, the fragments of level $1$ and the fragments of level $2$. As $L_{1,2}$ is trivial by definition, there exists an embedded sphere $\mathcal{S}_{0}$ where $L_{1,2}$ is embedded. Since embeddings of connected spatial graphs in a sphere are cellular, the resulting embedding $(L_{1,2},\mathcal{S}_{0})$ is cellular.
Now handles are attached to the upper hemisphere $\mathcal{S}_{u}$ of $\mathcal{S}_{0}$ to construct a surface on which $L_{1,2}\cup L_{3}$ is embedded in. For that, consider the set of disks $D_{u}$ that are the connected components of $\mathcal{S}_{u}\setminus\mathcal{G}$. Since each endpoint $p$ of a fragment in $L_{3}$ lies on the boundary $\partial D_{p}$ of at least one disk $D_{p}\in D_{u}$, an assignment $a(p)=D_{p}\in D_{u}$ with $p\in\partial D_{p}$ can be chosen.
@@
Denote by $F_{i}$ the set of fragments of $L_{3}$ with endpoints in $Cl_{i}$. Since $F_{i}$ is trivial, there is a sphere $S_{F_{i}}$ on which $F_{i}$ embeds in. For each set of endpoints $p_{k}$ of $F_{i}$ for which $a(p_{k})=D_{p_{k}}$, remove a disk $P_{p_{k}}$ from $S_{F_{i}}$ such that $p_{k}\in\partial P_{p_{k}}$ for all $k$ and $q\notin P_{p_{k}}$ for any endpoint $q\neq p_{k}$. This construction yields an $n_{i}$-punctured sphere $\mathcal{S}_{i}$ on which all fragments of $Cl_{i}$ are embedded in.
@@
Since the punctured spheres $\mathcal{S}_{i}$ are constructed to follow the non-conflicting fragments of level $3$, it is implied that the $\mathcal{S}_{i}$ are not intersecting. Removing the interiors of the disks that are in the image of $a$ from $\mathcal{S}_{u}$ allows to construct an oriented closed non-intersecting surface. This is done by identifying their boundaries with the boundaries of the punctures of $\mathcal{S}_{i}$ that respects the induced orientations on the boundaries of the surfaces.
@@
The punctured spheres on which $L_{4}$ embeds in do not intersect the surface where $L_{1,2}\cup L_{3}$ is embedded. This is due to there being four levels only: each level is attached to only one of the hemispheres and the crossing types between the levels are of only one type, making the attachment of handles to the upper hemisphere and to the lower hemisphere independent of one another. The result is an embedding of $\mathcal{G}=L_{1}\cup L_{2}\cup L_{3}\cup L_{4}$ in a handlebody of genus $\sum_{i}(n_{i}-1)+\sum_{j}(n_{j}-1)$.
It remains to be seen that the embedding is cellular. The embedding of $\mathcal{G}$ in the resulting handlebody is cellular since each $\mathcal{S}_{k}$ is attached to different disks of $\mathcal{S}_{0}\setminus\mathcal{G}$, and $F_{k}$ is cellular embedded in $\mathcal{S}_{k}$, and $L_{1,2}$ is cellular embedded in $\mathcal{S}_{0}$.
@@
The restriction to four levels in proposition 1 assures both the cellularity of the resulting embedding and the absence of self-intersections of the resulting surface. The construction of the proof of proposition 1 ensures neither property for a leveled embedding with five or more levels, as can be seen in the following example.
@@
Consider the leveled embedding $\mathcal{G}$ shown left in fig. 13, that is given by the spine list
$(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},3^{1},3^{2},3^{3},4^{1},4^{2},4^{3},5^{1},5^{2},5^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},3^{3},3^{2},3^{1},4^{3},4^{2},4^{1},5^{3},5^{2},5^{1}).$
We can apply the algorithm in the proof of proposition 1 to embed any four levels in a surface $\mathcal{S}$; without loss of generality choose the levels 1, 2, 4, 5. This leaves level 3 to be embedded. The endpoints of fragments of level 3 lie on the boundary of two disks of $\mathcal{S}\setminus\{L_{1}\cup L_{2}\cup L_{4}\cup L_{5}\}$. Denote by $D_{u}$ (respectively $D_{l}$) the disk which has interior points in common with the upper (resp. lower) hemisphere $S_{u}$ (resp. $S_{l}$) of the sphere that is produced in the first step of constructing $\mathcal{S}$. Extending $\mathcal{S}$ to accommodate $L_{3}$ cellularly requires a single cylinder $P$ to be placed under fragments $3^{1},3^{2},3^{3}$ such that $P$ is attached to $\mathcal{S}$ with one end glued to $D_{u}$ and the other to $D_{l}$. However, the resulting surface is non-orientable. This can be seen as follows: the order in which the three fragments are attached to the spine defines an orientation of the spine. As the three fragments do not conflict in $\mathcal{G}$ with respect to the spine, and $P$ is glued to $D_{u}$ on one end and to $D_{l}$ on the other, the two ends of $P$ are oriented with the same orientation as the spine. This is, if the orientation on one boundary component of the cylinder is clockwise, it is counterclockwise for the other. The boundary orientation does therefore not correspond to the orientation that is obtained when considering the two ends of the cylinder as naturally identified. This implies that the surface obtained by gluing $P$ to the oriented surface $\mathcal{S}$ is non-orientable.

### v2

@@ Abstract
For a given spatial graph $\mathcal{G}\subset\mathbb{R}^{3}$, we would like to find a closed orientable surface $\mathcal{S}$ embedded in $\mathbb{R}^{3}$ in which $\mathcal{G}$ is cellular embedded. However, for general $\mathcal{G}$ this is not possible. We therefore define a property of spatial graphs, called leveled, to show that for leveled spatial graphs with a small number of levels, a surface $\mathcal{S}$ can always be found. The argument is based on decomposing $\mathcal{G}$ into spatial subgraphs that can be placed on a sphere and on cylinders attached as handles, in such a way that the resulting surface contains a cellular embedding of $\mathcal{G}$.
We generalize the procedure to an algorithm that, if successful, constructs $\mathcal{S}$ for leveled spatial graphs with any number of levels. We conjecture that all connected leveled embeddings can be cellular embedded with the presented algorithm.
@@ 1 Introduction
Cellular embeddings, as a natural generalization of planar graphs, have in particular been studied in topological graph theory and combinatorics. They provide a combinatorial description of graph embeddings in surfaces Sta (78), and they are at the core of long-standing open conjectures such as the circular (or strong) embedding conjecture Jae (85).
@@
Given that graphs on surfaces have mainly be studied in topological graph theory that, in contrast to knot theory, traditionally considers all possible embeddings of an abstract graph, the scenario involving specific realizations of graphs within three-dimensional space, where surfaces must not intersect, remains underexplored. This is despite the setting being most relevant for applications taking place in the 3-dimensional Euclidean world. For example, building a graph-like structure using self-assembly of DNA molecules by origami folding requires first finding an unknotted A-trail through the desired spatial graph EMPS+ (17), MAG+ (20). When the graph is eulerian and cellular embedded in a surface, an unknotted A-trail can be found by following the boundary of the disks EMPS+ (17).
@@
Since we do not require the surface to be standardly embedded (e.g. a torus is allowed to be knotted), it is in principle possible to obtain all cellular embeddings of spatial graphs by starting from cellular embeddings of free bouquets of circles, on which subdivisions of edges and splittings of vertices are performed on the surface. However, our goal is to construct a surface for a given spatial graph. Note that rotation systems, although equivalent to cellular embeddings for abstract graphs Sta (78), do not ensure the described surface to be free of self-intersections.
@@
A natural approach to construct a surface for a given spatial graph is to glue together non-intersecting properly embedded disks that are bounded by some cycles of the graph. A sufficient albeit unnecessarily strong condition for the existence of such disks is $k$-flatness of the spatial graph NT (04). Indeed, it suffices to find a collection of cycles that cover all edges of the graph such that each edge is covered at most twice, and that bound disks satisfying the property above. The union of the disks and the spatial graph produces a cellular embedding of the spatial graph in the resulting surface. In case that every edge belongs to exactly two different cycles, the embedding is a circular (or strong) embedding. However, finding a collection of such disks is in general an intractable problem, as it includes detecting the unknot, non-trivial links with any number of components, Borromean links, and more complicated entanglements such as nonfree knot- and linkless spatial graphs such as the Kinoshita theta-curve.
@@
We therefore follow another approach to construct the surface. Without loss of generality, we can restrict to connected spatial graphs. Indeed, if distinct connected components are nontrivially linked, the spatial graph cannot be cellular embedded, and unlinked components can be considered independently. The idea is to start with a subgraph of the spatial graph that cellular embeds in the sphere, and then iteratively attach handles on which the remaining parts of the graph sit on. If the graph follows a longitude and a meridian of every handle, the resulting embedding is cellular. Since in our construction a longitude of the cylinder that generates a new handle is always followed by some subgraph, cellularity is implied if the two boundary components of the cylinder are attached to different disks of the already generated surface. If the two boundary components of the cylinder are attached to the same disk, it must be ensured that the graph also runs through the meridian of the resulting handle. Furthermore, conditions need to be found that guarantee that the construction results in an embedded surface, i.e., handles should be attached in such a way that they do not intersect the already constructed surface parts. For such a procedure to work, we require the spatial graph to have some properties and we call such spatial graphs leveled. They can be viewed as a generalization of planar graphs and are distinct from but related to open book embeddings BB (24).
@@ 2 Preliminaries
All graphs $G$ are undirected finite graphs. Graphs are allowed to have loops and parallel edges. The fragments (or bridges) of a graph $G$ with respect to a cycle $C$ of $G$ are the closures of the connected components of GCG-C Tut (56). Two fragments are conflicting if they either have pairs of endpoints on $C$ that alternate, or have at least three endpoints in common Wes (96). The conflict graph of a graph $G$ with respect to a cycle $C$ is the graph that has a vertex for each fragment of $G$ with respect to $C$, and two vertices are adjacent if and only if the corresponding fragments conflict. A connected graph is hamiltonian if it contains a cycle that visits every vertex exactly once, i.e., all its fragments with respect to the cycle are the closures of edges or loops. A graph embedding is an embedding $f:G\rightarrow S^{3}$ of a graph $G$ in $S^{3}$ up to ambient isotopy. The corresponding spatial graph $\mathcal{G}$ is the image of this embedding. A spatial graph is trivial if it embeds in $S^{2}$. An abstract graph is planar if it admits a trivial embedding. A diagram of a spatial graph $\mathcal{G}$ is the image of a projection of $\mathcal{G}$ onto $\mathbb{R}^{2}$ such that all intersections in the image are transversal, no vertices and at most two points are mapped to a crossing, and each double point is assigned as an over- or under-crossing. We focus on a special class of spatial graphs, the leveled graphs.
@@
We define an ordering of the fragments of $\mathcal{G}$ as follows: fix a diagram $D$ of a leveled spatial graph $\mathcal{G}$ such that the spine $C$ of $\mathcal{G}$ does not have crossings on $D$. If the fragment $f$ crosses over fragment $g$, then $f>g$. Fragments that are not comparable by the partial order are said to be on the same level. Note that conflicting fragments must be on different levels. Also, it is not possible that two fragments $f,g$ are such that $f>g$ and $g>f$. Indeed, if this was the case, the interior $\mathring{D_{f}}$ of $D_{f}$ could not be disjoint from $D_{g}$.
This definition allows us to assign the fragments of a leveled spatial graph to their level. The fragments that do not cross over any other fragments are at level $1$, denoted $L_{1}$.
@@
If the leveled spatial graph $\mathcal{G}$ has $n$ levels $L_{i},1\leq i\leq n$, there is an action of the cyclic group $C_{n}=\langle x\mid x^{n}=1\rangle$ on the partition of levels $\mathcal{L}=\{L_{1},\dots,L_{n}\}$. The action is given by sending $x\cdot L_{i}$ to $L_{i+1(\mathrm{mod}\ n)}.$
We say that a leveled spatial graph $\mathcal{G}$ is in representative leveling if for each fragment $f$ at level $i>1$, there exists a fragment $g$ at level i1i-1 such that $f$ and $g$ conflict. In other words, $\mathcal{G}$ is in representative leveling if every fragment belongs to the lowest possible level. Every leveled spatial graph can be put into representative leveling by assigning fragments to the lowest possible level. Note that the cyclic reordering of levels does not maintain representative leveling.
Every connected trivial spatial graph is a leveled graph with exactly one level. Consequently, all planar graphs admit a leveled embedding, but that is not true for all graphs. For example, the graph $K_{5}\_K_{5}$ that is obtained by joining two copies of $K_{5}$ by an edge (Figure 13) cannot be embedded as a leveled graph, since every choice of a cycle implies a non-planar fragment. Therefore, no spine can be found for $K_{5}\_K_{5}$.
@@
If the two fragments that have the vertex $v$ in common are non-conflicting, following $C$ starting from $v$ runs through all endpoints of one fragment before running through all endpoints of the other. Splitting $v$ as described above does not introduce an alternation of the endpoints, and therefore keeps the fragments non-conflicting. ∎
By repeatedly splitting vertices on the spine of a leveled spatial graph in the way described in Lemma 1, their degree can be reduced to three.
A hamiltonian leveled spatial graph $\mathcal{H}$ is a leveled spatial graph whose spine is a hamiltonian cycle of the underlying abstract graph. For constructing a cellular embedding for a hamiltonian leveled spatial graph, due to Lemma 1, it is possible to restrict without loss of generality to simple graphs, i.e. graphs without loops and parallel edges. In this case, all fragments are edges with two distinct endpoints. Hamiltonian leveled spatial graphs can be completely described combinatorially by their spine list that encodes the order in which fragments are attached to the spine together with their level. If several fragments share an endpoint, the order in the spine list is chosen such that the alternation of fragment labels is avoided. The spine list is unique up to permutation of the labels of parallel edges and cyclic permutations of its symbols, which corresponds to different choices of start points on the spine, see Figure 1.
@@ 3 Cellular embeddings of leveled spatial graphs with few levels
Assume $\mathcal{G}$ is in representative leveling. Let $L_{i}$ be the union of all fragments of level $i$, and let $L_{1,2}:=C\cup L_{1}\cup L_{2}$ be the subgraph of $\mathcal{G}$ that consists of the spine $C$ of $\mathcal{G}$, the fragments of level 11 and the fragments of level $2$. As $L_{1,2}$ is trivial by definition, there exists an embedded sphere $\mathcal{S}_{0}$ where $L_{1,2}$ is embedded. Since embeddings of connected spatial graphs in a sphere are cellular, the resulting embedding $(L_{1,2},\mathcal{S}_{0})$ is cellular.
Now handles are attached to the upper hemisphere $\mathcal{S}_{u}$ of $\mathcal{S}_{0}$ to construct a surface where $L_{1,2}\cup L_{3}$ is embedded. For that, consider the set of disks $D_{u}$ that are the connected components of $\mathcal{S}_{u}\setminus\mathcal{G}$. Since each endpoint $p$ of a fragment in $L_{3}$ lies on the boundary $\partial D_{p}$ of at least one disk $D_{p}\in D_{u}$, an assignment $a(p)=D_{p}\in D_{u}$ with $p\in\partial D_{p}$ can be chosen.
@@
Denote by $F_{i}$ the set of fragments of $L_{3}$ with endpoints in $Cl_{i}$. Since $F_{i}$ is trivial, there is a sphere $S_{F_{i}}$ where $F_{i}$ embeds. For each set of endpoints $p_{k}$ of $F_{i}$ for which $a(p_{k})=D_{p_{k}}$, remove a disk $P_{p_{k}}$ from $S_{F_{i}}$ such that $p_{k}\in\partial P_{p_{k}}$ for all $k$ and $q\notin P_{p_{k}}$ for any endpoint $q\neq p_{k}$. This construction yields an $n_{i}$-punctured sphere $\mathcal{S}_{i}$ where all fragments of $Cl_{i}$ are embedded.
@@
Since the punctured spheres $\mathcal{S}_{i}$ are constructed to follow the non-conflicting fragments of level $3$, it is implied that the $\mathcal{S}_{i}$ are not intersecting. Removing the interiors of the disks that are in the image of $a$ from $\mathcal{S}_{u}$ allows us to construct an oriented closed non-intersecting surface. That is done by identifying their boundaries with the boundaries of the punctures of $\mathcal{S}_{i}$ that respect the induced orientations on the boundaries of the surfaces.
@@
The punctured spheres where $L_{4}$ embeds do not intersect the surface where $L_{1,2}\cup L_{3}$ is embedded. That is due to there being four levels only: each level is attached to only one of the hemispheres and the crossing types between the levels are of only one type, making the attachment of handles to the upper hemisphere and to the lower hemisphere independent of one another. The result is an embedding of $\mathcal{G}=L_{1}\cup L_{2}\cup L_{3}\cup L_{4}$ in a surface of genus $\sum_{i}(n_{i}-1)+\sum_{j}(n_{j}-1)$.
It remains to be seen that the embedding is cellular. The embedding of $\mathcal{G}$ in the resulting surface is cellular since each $\mathcal{S}_{k}$ is attached to different disks of $\mathcal{S}_{0}\setminus\mathcal{G}$, and $F_{k}$ is cellular embedded in $\mathcal{S}_{k}$, and $L_{1,2}$ is cellular embedded in $\mathcal{S}_{0}$.
@@
The restriction to four levels in Proposition 1 assures both the absence of self-intersections of the resulting surface and the cellularity of the resulting embedding. The construction of the proof of Proposition 1 ensures neither property for a leveled embedding with five or more levels, as can be seen in the following example.
@@
Consider the leveled embedding $\mathcal{G}$ shown left in Figure 15, that is given by the spine list $(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},3^{1},3^{2},3^{3},$ $4^{1},4^{2},4^{3},5^{1},5^{2},5^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},$ $3^{3},3^{2},3^{1},4^{3},4^{2},4^{1},5^{3},5^{2},5^{1}).$ We can apply the algorithm in the proof of Proposition 1 to embed any four levels in a surface $\mathcal{S}$; without loss of generality choose the levels 1, 2, 4, 5. This leaves level 3 to be embedded. The endpoints of fragments of level 3 lie on the boundary of two disks of $\mathcal{S}\setminus\{L_{1}\cup L_{2}\cup L_{4}\cup L_{5}\}$. Denote by $D_{u}$ (respectively $D_{l}$) the disk which has interior points in common with the upper (resp. lower) hemisphere $S_{u}$ (resp. $S_{l}$) of the sphere that is produced in the first step of constructing $\mathcal{S}$. Extending $\mathcal{S}$ to accommodate $L_{3}$ cellularly requires a single cylinder $P$ to be placed under fragments $3^{1},3^{2},3^{3}$ such that $P$ is attached to $\mathcal{S}$ with one end glued to $D_{u}$ and the other to $D_{l}$. However, the resulting surface is non-orientable. This can be seen as follows: the order in which the three fragments are attached to the spine defines an orientation of the spine. As the three fragments do not conflict in $\mathcal{G}$ with respect to the spine, and $P$ is glued to $D_{u}$ on one end and to $D_{l}$ on the other, the two ends of $P$ are oriented with the same orientation as the spine. That is, if the orientation on one boundary component of the cylinder is clockwise, it is counterclockwise for the other. The boundary orientation does therefore not correspond to the orientation that is obtained when considering the two ends of the cylinder as naturally identified. This implies that the surface obtained by gluing $P$ to the oriented surface $\mathcal{S}$ is non-orientable.